OpenStudy (anonymous):

verify the functions are inverses: f(n)=-16+n/4 g(n)=4n+16 can anyone explain how I would go about answering this problem step-by-step

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

Compute f(g(n)) and g(f(n)) and verify that they are equal to n

OpenStudy (anonymous):

\[ f(g(n))=\frac{1}{4} (4 n+16)-4=n \]

OpenStudy (anonymous):

\[ g(f(n))=4 \left(\frac{n}{4}-4\right)+ 16=n \]

OpenStudy (anonymous):

@eliassaab thank you soooo much can you help me with one more? f(x)=x^2-9 g(x)=x+3

OpenStudy (anonymous):

@mathstudent55 help?????

OpenStudy (anonymous):

When you need to ask another problem, than post it as a new problem.